OpenStudy (anonymous):

L is the midpoint of CD. If CL=1/3x+8 and LD2/3x-4, find the length of CD.

OpenStudy (anonymous):

@Mertsj

OpenStudy (nikato):

First, u want to find x. Since L is the midpoint, those 2 segments are =. 1/3x+8=2/3x-4

OpenStudy (mertsj):

Solve the proportion nikato posted.

OpenStudy (anonymous):

12=1/3x is this correct?

OpenStudy (mertsj):

Cross multiply: 6x+16=3x-4

OpenStudy (nikato):

Correct. And now multiply both sides by 3 so x=36. Substitute this value for both and add them together

OpenStudy (anonymous):

@nikato i dont quite understand what you just asked me to do sorry

OpenStudy (nikato):

Ok. So we have x=36. Now substitute this value to both those expression to find the lengths of CL and LD

OpenStudy (anonymous):

@Mertsj 6.7=x is what i got from what you told me to solve

OpenStudy (mertsj):

The statement of the problem is not clear. Is this it: (1/3)x +8=(2/3)x-4 ???

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Then, adding 4 to both sides we have: (1/3)x+12=(2/3)x

OpenStudy (mertsj):

Subtracting (1/3)x from both sides we have: 12=(1/3)x

OpenStudy (mertsj):

Multiplying both sides by 3 we have: 36=x

OpenStudy (anonymous):

i think im following

OpenStudy (mertsj):

If x = 36 then (1/3)x+8=(1/3)(36)+8=12+8=20

OpenStudy (mertsj):

So half of the segment is 20 units and the entire segment is 40

OpenStudy (anonymous):

ok cool :)

OpenStudy (anonymous):

so CD =40?

OpenStudy (mertsj):

So half of the segment is 20 units and the entire segment is 40

OpenStudy (anonymous):

:) thanks